OpenStudy (anonymous):

factor and solve cot^2xcos^2x-cot^2x=0

OpenStudy (solomonzelman):

\(\Large\color{black}{\cot^2x \cos^2x-\cot^2x=0}\) So your first step, is to factor out of \(\Large\color{black}{\cot^2x }\), can you do that?

OpenStudy (redheadangel):

Here is another way to verify by working on both side together, cot^2x - cos^2x = cot^2x * cos^2x, where cot^2x = (cos^2x)/(sin^2x) Substitute cot^2x and multiply the second term with (sin^2x)/(sin^2x) for common denominator (cos^2x)/(sin^2x) -( cos^2x)(sin^2x)/(sin^2x) = (cos^2x)/(sin^2x) * cos^2x Multiply by sin^2x on both sides to eliminate denominator and factor, then cos^2x(1 - sin^2x) = cos^2x * cos^2x Devide by cos^2x on both sides 1- sin^2x = cos^2x, where 1 - sin^2x = cos^2x Then, cos^2x = cos^2x, both sides same.

OpenStudy (anonymous):

Yes, I factored it out

OpenStudy (anonymous):

Okay, so I factored it to \[Cot ^{2}\alpha(\cos ^{2}\alpha)(-1)=0\]

OpenStudy (solomonzelman):

\(\Large\color{black}{ \cot^2x \cos^2x - \cot^2x=0}\) \(\Large\color{black}{ \cot^2x (\cos^2x - 1)=0}\) Good? now apply the zero product property.

OpenStudy (anonymous):

What's the zero product property?